OpenStudy (anonymous):

If six out of 15 customers order a chicken sandwich, how many chicken sandwiches should a restaurant predict they will need to serve 270 people?

OpenStudy (anonymous):

108

OpenStudy (anonymous):

Is that a choice?

OpenStudy (warriorz13):

108

OpenStudy (anonymous):

6 out of 15 is .4% so multiply .4 by the 270 people. That leaves you with 108 chicken sandwhiches.

OpenStudy (anonymous):

yes thanks

OpenStudy (warriorz13):

270/15=18*6=108

OpenStudy (anonymous):

Your welcome,and thank you.